OpenStudy (anonymous):

Evaluate each polynomial for x = 3 and for x = −3. −2x^3 − 3x^2 + 4x + 2

OpenStudy (mertsj):

Can you plug in 3 for x and evaluate?

OpenStudy (anonymous):

ok so would it be -2*3^3-3*3^2+4*3+2

OpenStudy (mertsj):

\[-2(3)^3-3(3)^2+4(3)+2\]

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

\[-54-27+12+2=-67\]

OpenStudy (mertsj):

Now plug in -3

OpenStudy (anonymous):

2*-3^3-3*-3^2+4*-3+2

OpenStudy (mertsj):

\[-2(-3)^3-3(-3)^2+4(-3)+2=54-27-12+2=17\]

OpenStudy (anonymous):

ok thats the whole thing

OpenStudy (mertsj):

Yep

OpenStudy (anonymous):

thank you so much

OpenStudy (mertsj):

yw